The length of each side of the ABCD square is 2m. Charge at point A is + 6C and charge at point B is + 8C. How much work is required to take two electrons from point C to point D?


1
Expert's answer
2021-10-05T15:44:10-0400

The work required to move from point C to point D is "A = q(\varphi_C - \varphi_D)", where here "q = 2 e" is the charge and "\varphi_C, \varphi_D" are potentials due to electric field created by charges A, B at points C, D respectively. Let us denote the length of the side of the square "l = 2 m".

Potential due to point charge is "\varphi = \frac{k q }{r}", ("k = 9 \cdot 9^{10} \frac{N m^2}{C^2}" is Coulomb constant), and it is additive, hence "\varphi_C = k \left( \frac{q_1}{l} + \frac{q_2}{\sqrt{2} l} \right)", "\varphi_D = k \left( \frac{q_1}{\sqrt{2} l} + \frac{q_2}{l} \right)".

Therefore, the work is "A = 2 e (\varphi_C - \varphi_D) = 2 e k\left[ q_1\left( \frac{1}{l} - \frac{1}{\sqrt{2} l}\right) + q_2\left( \frac{1}{\sqrt{2} l} - \frac{1}{l}\right)\right] \approx 8.43 \cdot 10^{-10} J"